OpenStudy (anonymous):

You are making poster board boundary markers in the shape of square pyramids for a relay race. If you are including a ten -inch square as the bottom and you want the slant height to be twelve inches, how much poster board do you need?

OpenStudy (anonymous):

a.340 b.360 c.384 d.400 plzzz help...

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

someboody pleaze help

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

hero (hero):

Do you know the formula for square pyramid?

OpenStudy (anonymous):

V=a2h 3

OpenStudy (anonymous):

V = a^2 h/3

OpenStudy (anonymous):

is that it

OpenStudy (anonymous):

@Hero ???

hero (hero):

The slant height is different from the height going from the base to the top of the pyramid: http://www.ck12.org/geometry/Surface-Area-and-Volume-of-Pyramids/lesson/Pyramids/

OpenStudy (aum):

This is not a volume problem but a surface area problem. Find the area of 1 triangle that forms the side of the pyramid. Multiply it by 4. Add the area of the square base.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what would the answer be

OpenStudy (aum):

Area of triangle = 1/2 * base * height = 1/2 * 10 * 12. Multiply by 4. Area of base = 10 * 10 Add them up.

OpenStudy (anonymous):

oooh now i understand :)

OpenStudy (anonymous):

i got 340

OpenStudy (anonymous):

is that correct??

OpenStudy (aum):

I believe so.

OpenStudy (anonymous):

:) thanks

OpenStudy (aum):

yw.

OpenStudy (anonymous):

yaaaaaas

OpenStudy (anonymous):

i love you tanks

OpenStudy (yanasidlinskiy):

By the way, @Mr_Meagee \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)